A weight of 20 lbs is on an inclined plane making 30 degrees above the horizontal. What component of the weight causes the body to slide along the plane? What is the normal reaction of the plane?


1
Expert's answer
2021-01-29T16:55:39-0500

The horizontal component of the weight causes body to slide along the plane.

The normal reaction of the plane can be calculated as follows:


"N=mgcos\theta=20\ lbs\cdot32\ \dfrac{ft}{s^2}\cdot cos30^{\circ}=554\ \dfrac{ft\cdot lb}{s^2}."